Calculate the pH at 0, 25.0, 50.0, 75.0, 100, and 140% titration in the titration of both protons of the diprotic acid H2SO3 with 0.100M NaOH, starting with 25.0 mL of 1.00 M H2A.

Ka1=1.20*10^-2

Ka2=6.60*10^-8
Answer #1

I think it is H2SO4

Following reation occr

H2SO4----->H++HSO4-

HSO4-------->H++SO42-

1.

NaOH=0

Ka1=[H+][HSO4-]/[H2SO4]=x*x/1=1.20*10-2

x=0.1095

[H+]=0.1095

pH=-log[H+]=0.960

2. NaOH=25%

1st buffer region, both H2A and OH- present

Ka1 > >Ka2, second equilibrium makes little contribution, pH calculated like a normal buffer solution

So half way to equivalence

[H2SO4]=[OH-]

So

[H+]=Ka1=1.20*10-2

pH=1.92

3.

50%

solution is like that of a salt of a diprotic acid

[H+]=((Ka2*[NaOH]+Kw)/(1+[NaOH]/Ka1))0.5

Kw=10-14

[H+]=2.66*10-5

pH=4.575

(4)

75%

pH in 2nd buffer region

2nd buffer region, both HA- and A2- present if Ka1 > 103 Ka2, first equilibrium makes little contribution, pH calculated like a normal buffer solution half way to equivalence,[HSO3-]=[SO42-]

[H+]=Ka2=6.60*10-8

pH=7.18

(5)

100%

pH at the second equivalence point Like a salt of A2-, main equilibrium is

SO42-+H2O---->OH-+HSO4-

Kb1=Kw/Ka2=10-14/6.60*10-8=1.515*10-7

[OH-]=(Kb1*1)0.5=3.892*10-4

pOH=3.40977

pH=70.59

(6)

140%

pH beyond 2nd equiv. pt.

treated like the addition of strong base to water.

pH=pw=14